# General Order Virtual Element Methods for Neumann Boundary Optimal Control Problems in Saddle Point Formulation

### v1

@@ 4. Error Estimates
Using the above estimates in (38) and simplifying $\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}$, we get (36).
@@
Let $(\boldsymbol{x},p)=((y,u),p)\in\left(\mathrm{H}^{s+1}\!(\Omega)\cap\mathrm{H}^{1}_{0}\!(\Omega)\right)\times\mathrm{H}^{r}\!(\Gamma_{C})\times\left(\mathrm{H}^{s+1}\!(\Omega)\cap\mathrm{H}^{1}_{0}\!(\Omega)\right)$ be the solution to Problem 3, with $0\leq s\leq k$ and $10\leq r\leq k+$ 1. Let $f\in\mathrm{H}^{t}\!(\Omega)$ and $y_{\mathrm{d}}\in\mathrm{H}^{\tau}\!(\Omega_{\text{obs}})$ be the right-hand side of Problem 3, with $10\leq t\leq k+$ 1 and $10\leq\tau\leq k+$ 1. Then, there exists a constant $C$ independent of $h$ such that the unique solution $((y_{h},u_{h}),p_{h})\in X_{h}\times Y_{h}$ to Problem 4 satisfies the following error estimate:
(40) $\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}+\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}&\leq C\big{(}h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\\ &\qquad\qquad\qquad+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\big{)}\end{split}$
@@
(41) $\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}\leq\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}+\left\lVert{\boldsymbol{x}_{I}-\boldsymbol{x}_{h}}\right\rVert_{X}\,,$
@@
(42) $\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}&=\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}+\left\lVert{u-u_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\\ &\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\end{split}$
@@
(43) $\alpha_{\ast}\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}^{2}$ $\leq{\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$
@@
(44) ${\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$
@@
${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)$ $={m}_{h}\left(y,y_{{\mathcal{B}}_{h}}-y_{h}\right)-\left(y,y_{{\mathcal{B}}_{h}}-y_{h}\right)_{\Omega_{\text{obs}}}$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\qquad\qquad\qquad\qquad-\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y,(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(\Pi^{0,E}_{k}y-y,(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$ $={\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{B}}\left(p,\boldsymbol{\varepsilon}_{h}\right)$
$\quad\quad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
$\qquad\qquad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
$\qquad\qquad\qquad+\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
(45) ${a}_{h}\left(y_{{\mathcal{B}}_{h}}-y_{h},p_{I}\right)-{a}\left(y_{{\mathcal{B}}_{h}}-y_{h},p\right)\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert y_{{\mathcal{B}}_{h}}-y_{h}\right\rvert\right\rvert\right\rvert_{Y}\,,$
following the same steps applied in the Proof of Lemma 5 (see (39)). The term $\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}$ is estimated using the Cauchy-Schwarz inequality and (33), obtaining
(46) $\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}\lesssim h^{s+1}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lVert{u_{I}-u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\,.$
@@
(47) $\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\qquad\qquad=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
(48) ${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)\lesssim\left(h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\right)\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$
@@
(49) $\begin{split}\left\lVert{\boldsymbol{x}_{I}-\boldsymbol{x}_{h}}\right\rVert_{X}&\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\\ &\qquad\qquad\qquad+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,.\end{split}$
@@
(50) $\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}&\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\\ &\qquad\qquad\qquad+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,.\end{split}$
Now, let us focus on the estimate of $\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}$. Let $p_{I}\in Y_{h}$, interpolating $p$, be given by Lemma 3. Applying the triangle inequality, we have
(51) $\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\left\lvert\left\lvert\left\lvert p-p_{I}\right\rvert\right\rvert\right\rvert_{Y}+\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}$
From (33), we obtain that
(52) $\left\lvert\left\lvert\left\lvert p-p_{I}\right\rvert\right\rvert\right\rvert_{Y}\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\,.$
@@
(53) $\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\frac{1}{\beta_{\ast}}\sup_{\boldsymbol{\varsigma}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)}{\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}}\,.$
@@
(54) ${\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)$ $={\mathcal{B}}_{h}\left(p_{I},\boldsymbol{\varsigma}_{h}\right)\mp{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)$
@@
The first two terms are estimated as
(55) ${a}_{h}\left(w_{h},p_{I}\right)-{a}\left(w_{h},p\right)\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}$
following, as previously, the same steps applied in the Proof of Lemma 5 (see (39)). Then, applying the Cauchy-Schwarz inequality, a trace inequality together with a Poincarè inequality and (33), the term $\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}$ in (54) is bounded as
@@
(56) $\begin{split}\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}&\leq\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{p-p_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\\ &\lesssim\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{\nabla(p-p_{I})}\right\rVert_{\mathrm{L}^{2}\!(E)}\\ &\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\,.\end{split}$
The last two terms of (54), applying Problems (3) and (4), become
@@
(57) $\begin{split}{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)&=-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)+\left(y_{\mathrm{d}},w_{h}\right)_{\Omega_{\text{obs}}}+{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)\\ &\qquad\qquad\qquad\qquad-\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},\Pi^{0,E}_{k}w_{h}\right)_{E}\,,\end{split}$
where applying the Cauchy-Schwarz inequality and (31), we get
(58) $\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}},w_{h}-\Pi^{0,E}_{k}w_{h}\right)_{E}$
$\qquad\qquad\qquad=\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},w_{h}-\Pi^{0,E}_{k}w_{h}\right)_{E}$
$\qquad\qquad\qquad\leq\sum_{E\in\mathcal{M}_{h}\cap\Omega_{\text{obs}}}\left\lVert{y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{w_{h}-\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}$
@@
$\qquad\qquad\qquad\lesssim h^{\tau}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}h\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\lesssim h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$

### v2

@@ 4 Error Estimates
Using the above estimates in (47) and simplifying $\left\lvert\left\lvert\left\lvert e_{h}\right\rvert\right\rvert\right\rvert_{Y}$, we get (45). ∎
@@
Let $(\boldsymbol{x},p)=((y,u),p)\in\left(\mathrm{H}^{s+1}\!(\Omega)\cap\mathrm{H}^{1}_{0}\!(\Omega)\right)\times\mathrm{H}^{r}\!(\Gamma_{C})\times\left(\mathrm{H}^{s+1}\!(\Omega)\cap\mathrm{H}^{1}_{0}\!(\Omega)\right)$ be the solution to Problem 3, with $0\leq s\leq k$ and $0\leq r\leq k+1$. Let $f\in\mathrm{H}^{t}\!(\Omega)$ and $y_{\mathrm{d}}\in\mathrm{H}^{\tau}\!(\Omega_{\text{obs}})$ be the right-hand side of Problem 3, with $0\leq t\leq k+1$ and $0\leq\tau\leq k+1$. Then, there exists a constant $C$ independent of $h$ such that the unique solution $((y_{h},u_{h}),p_{h})\in X_{h}\times Y_{h}$ to Problem 4 satisfies the following error estimate:
$\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}+\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}&\leq C\big(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\\ &\qquad\qquad\qquad+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\big)\end{split}$ (49)
@@
$\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}\leq\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}+\left\lVert{\boldsymbol{x}_{I}-\boldsymbol{x}_{h}}\right\rVert_{X}\,$, (50)
@@
$\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{I}}\right\rVert_{X}&=\left\lvert\left\lvert\left\lvert y-y_{{\mathcal{B}}_{h}}\right\rvert\right\rvert\right\rvert_{Y}+\left\lVert{u-u_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\\ &\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\end{split}$ (51)
@@
$\alpha_{\ast}\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}^{2}$ $\leq{\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$ (52)
@@
${\mathcal{A}}_{h}\left(\boldsymbol{x}_{I}-\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)\lesssim\left(h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\right)\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$ (53)
@@
${\mathcal{A}}_{h}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)$ $={m}_{h}\left(y,y_{{\mathcal{B}}_{h}}-y_{h}\right)-\left(y,y_{{\mathcal{B}}_{h}}-y_{h}\right)_{\Omega_{\text{obs}}}$ (54)
$=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(\Pi^{0,E}_{k}y,\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$\qquad\qquad\qquad\qquad-\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y,(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
$=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(\Pi^{0,E}_{k}y-y,(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)$ $={\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{B}}\left(p,\boldsymbol{\varepsilon}_{h}\right)$ (55)
$\quad\quad+\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
$\qquad\qquad+\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
$\qquad\qquad\qquad+\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
${a}_{h}\left(y_{{\mathcal{B}}_{h}}-y_{h},p_{I}\right)-{a}\left(y_{{\mathcal{B}}_{h}}-y_{h},p\right)\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert y_{{\mathcal{B}}_{h}}-y_{h}\right\rvert\right\rvert\right\rvert_{Y}\,$, (56)
following the same steps applied in the Proof of Lemma 5 (see (48)). The term $\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}$ is estimated using the Cauchy-Schwarz inequality, a discrete trace inequality together with a Poincaré inequality and (41), obtaining
$\left(u_{I}-u_{h},p-p_{I}\right)_{\Gamma_{C}}\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lVert{u_{I}-u_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\,.$ (57)
@@
$\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$ (58)
$\qquad\qquad=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},(y_{{\mathcal{B}}_{h}}-y_{h})-\Pi^{0,E}_{k}(y_{{\mathcal{B}}_{h}}-y_{h})\right)_{E}$
@@
${\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varepsilon}_{h}\right)-{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varepsilon}_{h}\right)\lesssim\left(h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\right)\left\lVert{\boldsymbol{\varepsilon}_{h}}\right\rVert_{X}\,.$ (59)
@@
$\begin{split}\left\lVert{\boldsymbol{x}_{I}-\boldsymbol{x}_{h}}\right\rVert_{X}&\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\\ &\qquad\qquad\qquad+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,.\end{split}$ (60)
@@
$\begin{split}\left\lVert{\boldsymbol{x}-\boldsymbol{x}_{h}}\right\rVert_{X}&\lesssim h^{s}\left\lvert{y}\right\rvert_{s+1,\Omega}+h^{r}\left\lvert{u}\right\rvert_{r,\Gamma_{C}}\\ &\qquad\qquad\qquad+h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+h^{t+1}\left\lvert{f}\right\rvert_{t,\Omega}+h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\,.\end{split}$ (61)
Now, let us focus on the estimate of $\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}$. Let $p_{I}\in Y_{h}$, interpolating $p$, be given by Lemma 3. Applying the triangle inequality and (42), we have
$\left\lvert\left\lvert\left\lvert p-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\left\lvert\left\lvert\left\lvert p-p_{I}\right\rvert\right\rvert\right\rvert_{Y}+\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}+\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$ (62)
@@
$\left\lvert\left\lvert\left\lvert p_{I}-p_{h}\right\rvert\right\rvert\right\rvert_{Y}\leq\frac{1}{\beta_{\ast}}\sup_{\boldsymbol{\varsigma}_{h}\in X_{h}}\frac{{\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)}{\left\lVert{\boldsymbol{\varsigma}_{h}}\right\rVert_{X}}\,.$ (63)
@@
${\mathcal{B}}_{h}\left(p_{I}-p_{h},\boldsymbol{\varsigma}_{h}\right)$ $={\mathcal{B}}_{h}\left(p_{I},\boldsymbol{\varsigma}_{h}\right)\mp{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)$ (64)
@@
The first two terms are estimated similarly as in Lemma 5 (see (48)), hence adding and subtracting $\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(w_{h},p\right)$, applying the continuity of ${a}_{h}$ (26) and the estimate (42), we get
${a}_{h}\left(w_{h},p_{I}\right)-{a}\left(w_{h},p\right)$ $=\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(w_{h},p_{I}-p\right)+{a}^{E}_{h}\left(w_{h},p\right)-{a}^{E}\left(w_{h},p\right)$ (65)
$\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}+\sum_{E\in\mathcal{M}_{h}}{a}^{E}_{h}\left(w_{h},p\right)-{a}^{E}\left(w_{h},p\right)$
$\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\,$,
where the last estimate can be derived for the diffusion and the reaction term, thanks to their symmetry, as in [5, Theorem 5]. For what concerns the transport term, we derive the estimate as in [23, Theorem 3]. Then, applying the Cauchy-Schwarz inequality, a discrete trace inequality together with a Poincaré inequality and (41), the term $\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}$ in (64) is bounded as
@@
$\begin{split}\left(v_{h},p-p_{I}\right)_{\Gamma_{C}}&\leq\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\left\lVert{p-p_{I}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\lesssim h^{s}\left\lvert{p}\right\rvert_{s+1,\Omega}\left\lVert{v_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Gamma_{C})}\,.\end{split}$ (66)
The last two terms of (64), applying Problems 3 and 4, become
@@
$\begin{split}{\mathcal{B}}\left(p,\boldsymbol{\varsigma}_{h}\right)-{\mathcal{B}}_{h}\left(p_{h},\boldsymbol{\varsigma}_{h}\right)&=-{\mathcal{A}}\left(\boldsymbol{x},\boldsymbol{\varsigma}_{h}\right)+\left(y_{\mathrm{d}},w_{h}\right)_{\Omega_{\text{obs}}}+{\mathcal{A}}_{h}\left(\boldsymbol{x}_{h},\boldsymbol{\varsigma}_{h}\right)\\ &\qquad\qquad\qquad\qquad-\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}},\Pi^{0,E}_{k}w_{h}\right)_{E}\,.\end{split}$ (67)
Applying the Cauchy-Schwarz inequality and (39), we get
@@
$\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}},w_{h}-\Pi^{0,E}_{k}w_{h}\right)_{E}=\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left(y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}},w_{h}-\Pi^{0,E}_{k}w_{h}\right)_{E}\\ \leq\sum_{E\in\mathcal{M}_{h,\text{obs}}}\left\lVert{y_{\mathrm{d}}-\Pi^{0,E}_{k}y_{\mathrm{d}}}\right\rVert_{\mathrm{L}^{2}\!(E)}\left\lVert{w_{h}-\Pi^{0,E}_{k}w_{h}}\right\rVert_{\mathrm{L}^{2}\!(E)}\\ \lesssim h^{\tau}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}h\left\lVert{\nabla w_{h}}\right\rVert_{\mathrm{L}^{2}\!(\Omega)}\lesssim h^{\tau+1}\left\lvert{y_{\mathrm{d}}}\right\rvert_{\tau,\Omega_{\text{obs}}}\left\lvert\left\lvert\left\lvert w_{h}\right\rvert\right\rvert\right\rvert_{Y}\,.$ (68)